LUCI1/AO Commissionng

01:51 Steve open telescope

02:00 Greg startup SX WFS. Doug arrived at LBTO

02:25 Set Pointing and collimate on on-axis star

02:33 Preset to off-axis guide star
  • HD82906
  • R,I 9.0 mag
  • H,K 9.1 mag

02:40 Preset to A45_R8_0030, ACE, GS=5, RS=0, PA=180

02:45 iif_preset failed
  • problems with modifications

02:50 run irc_preset, iic[0] from LBTO account
  • start bin1, change to bin 2
  • error in table_LUCIFER_ACE-AO.txt
  • fix it

02:56 irc_runao
  • communication error between AOS and AOArbitrator
  • AOARB.L |ERR| 5130|2016-04-22 02:59:18.142717| SERIALIZATION > Error in message deserialization from AOS.L (014000/2012)
  • Fri Apr 22 03:04:18.074 2016 aos.AcquireRef{FLAO}.failed - AcquireRef {FLAO}[15] - Command timeout - -13107: Arbitrator interface generic error

03:05 Loop closed and optical gain loop running

03:18 GoptGUI died, seg fault

03:19 Setting up LUCI1 for subwindows

03:44 Opt loop data
  • 034404

03:50 Offset to place star on LUCI
  • Skip Frame
  • Loop open

04:03 Error during Recover Failure
  • had big T/T when I renormalized gain and applied to Gopt
  • Loop opened
  • Shell rested and went into failure
  • During Recover Failure get the output eo
4:03:36.218360|             MAIN > BCUs 2-2 ready!
adsec.L            |INF|      2155|2016-04-22 04:03:36.218555|             MAIN > Ok.  Switch Bcu are ready
adsec.L            |INF|      2156|2016-04-22 04:03:36.218958|             MAIN > housekeeper.L test started.
adsec.L            |INF|      2157|2016-04-22 04:03:47.275426|             MAIN > housekeeper.L non yet started or giving wrong reading in the CHECK DEW POINT variable (Nan, Inf, >200 deg or <0).
20162204 04:04:12.3448 Alert: Warning: recoverFailure not executed (retry)
adsecarb.L         |WAR|      1450|2016-04-22 04:04:12.344025|              IDL > Command returns retryable error : IDL File error
adsecarb.L         |INF|      1451|2016-04-22 04:04:12.344431|   COMMANDHANDLER > Command not executed (status RETRY) (errstr IDL File error)
adsec.L            |INF|      2160|2016-04-22 04:04:12.335390|             MAIN > Reset for fault recovery failed.
Last message repeated
adsec.L            |INF|      2161|2016-04-22 04:04:12.335496|             MAIN > IDL File error
adsec.L            |ERR|      2162|2016-04-22 04:04:50.587424|      ADAM-MODBUS > [TcpReceiveException] Error receiving data (errno=25 ), socket closed. (code -15003) TcpConnection receive error File TcpConnection.cpp line 130
adsec.L            |INF|      2163|2016-04-22 04:04:50.587514|             MAIN > adamModbusReadOutput_wrap found TCP connection not initialized ...
adsec.L            |INF|      2164|2016-04-22 0

04:20 Alfio notice that when the loop is closed with StartAO there is a larg clover leaf mode applied.
  • Find that the acquireRef start in bin1 at one re-rotator angle then switch to bin 2 at another re-rotator angle. The StartAO was closing the loop before the re-rotator got to its new position
  • Alfio install a 10 second delay. However, very seldom need this. Will try to make the re-rotator change request earlier in script

05:41 Opt Loop data, bin 2
  • NCPA ON
    • 054219
    • 054302
    • Dark Taken
    • lower TT gain by 0.1 down to 0.47
    • 055016
    • Remove H filter. Now only FeII
    • new Dark
      • 055427
      • 055516
      • 055542
  • No NCPA
    • 055706
    • 055750
    • 055825

06:33 Bin 1, 6.5 (see 7.8)
  • 063227 (lowered ho1 and ho2 by 0.05)
  • 063325 ( back to original)
  • 063827
  • 063904 (lowered ho1 and ho2 by 0.05)
  • 064008 (lowered ho1 and ho2 by 0.1)

06:48 (50000 counts)
  • 064956 ( ho1, h02 down by 0.1)
  • 065101 ( ho1, h02 down by 0.05)
  • 065155
  • NCPA ON
    • 065311 ( down by 0.05)
    • 065552
    • 065733
  • NCPA OFF

08:22 Arcetri work on Optical Gain routines

08:23 Start collecting bin 1-4 performance data

08:26 Preset to AO510, GS=1, RS=0, PA=180
  • Bin 1
  • catalog 6.8
  • measured 8.4
  • Seeing 1.0
  • Rot angle = 32 (low NCPA)
  • New Dark
  • NCPA ON
    • 084117
    • 084146
    • 084215
  • NCPA OFF
    • 084252
    • 084318
    • 084348

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20160422_084117 ??? 8.5 291544 10 1.26 0.51 062006 1 300 990 0.4 0.2 0.2 3 237 ON 51.3 FeII clear 172 50 ONSKY zero.fits  
20160422_084146 ??? 8.5 291320 9 1.13 0.51 062006 1 300 990 0.4 0.2 0.2 3 243 ON 52.1 FeII clear 172 50 ONSKY zero.fits - long stop of 53 frames at frame 3750
20160422_084215 ??? 8.5 291089 8 1.20 0.55 062006 1 300 990 0.4 0.2 0.2 3 237 ON 47.6 FeII clear 172 50 ONSKY zero.fits - long stop of 67 frames at frame 1026
IDL string: ['20160422_084117','20160422_084146','20160422_084215']
SR 0 (NaN)
SEEING (WFS) 0.519861 (0.0222534)
SEEING (DIMM) 1.19669 (0.0658244)

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20160422_084252 ??? 8.5 290796 9 1.13 0.47 062006 1 300 990 0.4 0.2 0.2 3 233 ON 54.2 FeII clear 172 50 ONSKY zero.fits - long stop of 82 frames at frame 2280
20160422_084318 ??? 8.5 290585 9 0.99 0.50 062006 1 300 990 0.4 0.2 0.2 3 234 ON 54.6 FeII clear 172 50 ONSKY zero.fits  
20160422_084348 ??? 8.6 290334 9 1.00 0.53 062006 1 300 990 0.4 0.2 0.2 3 224 ON 52.9 FeII clear 172 50 ONSKY zero.fits - long stop of 52 frames at frame 1832
IDL string: ['20160422_084252','20160422_084318','20160422_084348']
SR 0 (NaN)
SEEING (WFS) 0.500996 (0.0275204)
SEEING (DIMM) 1.04332 (0.0764566)

08:44 Preset to AO510, GS=4, RS=0, PA=-30
  • Bin 1
  • catalog 6.8
  • measured 8.4
  • Seeing 1.0
  • Rot angle = 205 (high NCPA)
  • New Dark
  • NCPA ON
    • 085416
    • 085447
    • 085511
  • NCPA OFF
    • 085448
    • 085619 (a bit of stripping)
    • 085707

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20160422_085416 ??? 8.5 284778 11 1.07 0.57 062006 1 300 990 0.5 0.3 0.2 3 245 ON 53.3 FeII clear 172 50 ONSKY zero.fits  
20160422_085447 ??? 8.5 284486 12 0.94 0.56 062006 1 300 990 0.5 0.3 0.2 3 243 ON 52.0 FeII clear 172 50 ONSKY zero.fits - long stop of 55 frames at frame 1328 - Background correction failed
20160422_085511 ??? 8.5 284258 11 1.00 0.52 062006 1 300 990 0.5 0.3 0.2 3 239 ON 54.3 FeII clear 172 50 ONSKY zero.fits - long stop of 51 frames at frame 1477
IDL string: ['20160422_085416','20160422_085447','20160422_085511']
SR 0 (NaN)
SEEING (WFS) 0.547086 (0.0280114)
SEEING (DIMM) 1.00472 (0.0659954)

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20160422_085548 ??? 8.5 283906 11 1.08 0.56 062006 1 300 990 0.5 0.3 0.2 3 239 ON 51.8 FeII clear 172 50 ONSKY zero.fits - long stop of 51 frames at frame 3237
20160422_085619 ??? 8.5 283609 10 1.21 0.78 062006 1 300 990 0.5 0.3 0.2 3 244 ON 43.4 FeII clear 172 50 ONSKY zero.fits - long stop of 65 frames at frame 3923
20160422_085707 ??? 8.4 283149 10 1.19 0.56 062006 1 300 990 0.5 0.3 0.2 3 253 ON 46.8 FeII clear 172 50 ONSKY zero.fits - long stop of 64 frames at frame 3187 - Background correction failed
IDL string: ['20160422_085511','20160422_085548','20160422_085619','20160422_085707']
SR 0 (NaN)
SEEING (WFS) 0.605699 (0.11956)
SEEING (DIMM) 1.11929 (0.0984786)

09:08 Preset to AO608, GS=1, RS=0, PA=0
  • Ooops, Bin 1 again

09:20 Ooops, table_LUCIFER_ACE-AO.bad_seeing.txt was incorrect.
  • fixed

09:50 Preset to AO590, GS=6, RS=0, PA=0
  • Bin 2
  • catalog 9.8
  • measured 11.3
  • Seeing 1.3
  • *GETTING KO* occasionally on Gopt GUI
    • Could have been that the Gopt loop was not closed soon enough and there was some gain drift
  • Rot angle = -19 (low NCPA)
  • New Dark
  • NCPA ON
    • 100224
    • 100254
    • 100322
  • NCPA OFF
    • 100409
    • 100432
    • 100456

10:10 Preset to AO590, GS=4, RS=0, PA=180
  • Bin 2
  • catalog 9.8
  • measured 11.1
  • Seeing 1.0
  • Rot angle = 150 (high NCPA)
  • NCPA ON
    • 102133
    • 102159
    • 102222
  • NCPA OFF
    • 102314
    • 102340
    • 102404

10:44 Preset to AO671, GS=21, RS=0, PA=90
  • Bin 1
  • elevation 59
  • catalog 7.3
  • measured 8.0
  • Seeing 1.0
  • Rot angle = 203 (high NCPA)
  • New Dark
  • GOPT OFF
    • 105812
    • 105838
    • 105907
  • GOPT ON, NO NCPA
    • 110231
    • 110259
    • 110321
  • NCPA ON
    • 110426
    • 110452
    • 110516

11:13 Preset to AO818, GS=24, RS=0, PA=0
  • Bin 1
  • elevation 35
  • catalog 7.9
  • measured 8.6
  • Seeing 1.0
  • Rot angle = 185 (high NCPA)
  • New Dark
  • GOPT ON
    • 113421
    • 113448
    • 113434
  • NCPA ON
    • 113728
    • 113752

11:55 Preset to AO674, GS=22, RS=0, PA=60
  • Bin 1
  • elevation 74
  • catalog 8.3
  • measured 9.4
  • Seeing 0.9
  • Rot angle = 185 (high NCPA)
  • NCPA ON
    • 115921
    • 120222
    • 120044
    • 120306
    • 120431

12:15 Too bright to test bin3. Closing

-- %USERSIG{DougMiller - 2016-04-22}%
Topic revision: r5 - 09 May 2016, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback